6/11/99 day shift summary

Sasha Glazov (glazov@hep.uchicago.edu)
Fri, 11 Jun 1999 15:46:41 -0500

Shift Summary for hhh shift, dd/mmm/99

On Shift: R. Ray, S. Glazov

Hours of beam delivered at all intensities: 7 hours
Hours of beam delivered at nominal intensity: 5 hours
Hours of E832 data (special runs not included): 7 hours
Hours of special runs (laser scans, muon runs...): ----

Accounting of lost time:

No beam from Tevatron

========================================================================
Runs ended this shift:
Run 13524 - 3.0M events - BEAM_832
Run 13526 - 2.2M events - BEAM_832
Run 13527 - 1.0M events - BEAM_832

Tapes brought to FCC this shift:

None

========================================================================
Accesses during shift:

At 14.00 beam was stopped for at least 1 hour. Ed, Jason, Elizabeth and
Miguel went on access. Ended at 14.30

========================================================================
Overview of shift:
E832 datataking

The overall gain of CsI is driftting due to the change
of the temperature

DC illuminations show many out of time hits. This is
due to the beam structure but also could be due to increased
noise.
========================================================================
Description of Shift:

7:00 Run 13524 in beam 832 configuration in progress.
7:25 HCC gets full. Reset/Kfrend --- works.
7:50 Pipeline errors. Same cure.
8:00 Ed changes prescale for trigger 2 from 60 to 20 to
to speed up online calibration
(Warning! watch /ktev3a disk space !)
8:40 Start new run (13526)
8:50 Onsplit wants to write out laser data (tape QKLA02)
Worked hard for ~3 hours. Success. Tape is still in
the tape drive.
10:00 RR studies online histograms/finds problems/mails experts.
Problems seen in cosmic#summary, fpedcal#load, muon#summary,
pipe#summary, t3#summary, ytf#summary -- you might like to
see them as well (or look into the log book for details).

10:30 Ed and SG compare 97 vs 99 run for kaon yield and trigger
rate. It looks like we get >30% MORE kaons while the rate of
trigger elements (for regenerator, V1, MU2) is basically the same.
Ed called Rick K but he also does not know the reason of such
great progress.

11:00 Still problems with a block 3031 of CsI. (see previous shift reports)
Ed observes low electron occupancy. He thinks that voltage
for this channel is far off.

11:20 R.Tesarek calls --> he found problems with HA ADC channel 23.
(which has too low average). SG looked at the signals with
scope, they look normal.

11:30 Problem with beam for estimated 20 min.

11:56 Jim updates RC vs Onsplit checking tool. Also onsplit software
is now updated to ignore 0 length files (works fine).
Jim is almost ready to declare that onsplit cache structures
are available on online machines the same way as on kpasa/ksera.
This would simplify running of calibration jobs over many
100 MB long files.

11:55 Stop run 13526. HV on CsI channel 3031 is changed by Val.
12:10 Start run 13527.
12:17 Beam is back.
12:20 New bottles of argon-ethane are installed at east side.
13:20 Tevatron attempts to improve beam structure. Intensity
assumed to vary due to vacuum work in the linac.
As the result we don't have any beam.

13:50 No beam. Ed calls the control room and asks for access.
This is Oked --> at least 1 hour no beam.

14:00 Ed, Jason, Miguel and Elizabeth go on access:
Ed & Jason -- study channel 3031
Miguel -- Changes threshold of DC 1 to 50mV. The hope is
to see less extra hits on DC for the same intensity. Other
chambers stay at 40mV to monitor the beam conditions.
Elizabeth -- adjusts MA 621 output width to 15 ns (as for MA-LO)

14:40 Val changes the HV of the famous channel (3031) back.

15:25 Stop the cosmic run (13529). Start new one.